OpenStudy (please_help):

Please solve using the Scientific rule: 10 X 10^-3

OpenStudy (kira_yamato):

= 10 ^(1 - 3) = 10^-2

OpenStudy (please_help):

Thank you. :)

OpenStudy (anonymous):

10^ 1 X 10^-3= 10 ^-2

OpenStudy (please_help):

Thank you as well! Great group.

OpenStudy (anonymous):

What's the scientific rule?